True/False:

a) If the Consumer Price Index is 125% of the base year in 2006 and 135% of the base year in 2007, the CPI increased by 10% from 2006-2007

b) When Internal Rate of Return is greater than Cost Benefit Ratio we should make the investment

c) Demand curves for non-rival goods are summed vertically

d) For any given good, Willingness to Accept is often larger than Willingness to Pay

Please provide explanation for why its True or False
Answer #1

A. False

Explanation: The % increase in CPI = (135 - 125)//125 * 100 = 10/125 * 100 = 1000/125 = 8%

B. False

Explanation These two are not compared.

C. True

Explanation: Demand curves for rival goods are summed horizontally and for the non-rival goods it is summed vertically.

D. False

Explanation: For any good, willingness to pay is often higher.